OpenStudy (anonymous):

Solve two times square root of the quantity x+8 = three times the square root of the quantity x-2

OpenStudy (anonymous):

In mathematical terms: 2sqrt(x+8) = 3sqrt(x-2) By squaring both sides: (2)^2 [sqrt(x+8)]^2 = (3)^2 [sqrt(x-2)]^2 4 (x+8) = 9 (x-2) 4x + 32 = 9x - 18 32 + 18 = 9x - 4x x = ?

OpenStudy (anonymous):

thanks so much! @dauspex